40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Dec 17 12:32:35 2021, Koji, Update, SUS, Remaining task for 2021 
    Reply  Fri Dec 17 19:19:42 2021, Koji, Update, SUS, Remaining task for 2021 PXL_20211218_030145356.MP.jpg
       Reply  Sun Dec 19 07:52:51 2021, Anchal, Update, SUS, Remaining task for 2021 
    Reply  Wed Dec 22 18:16:23 2021, Koji, Update, SUS, Remaining task for 2021 ITMX_2016_07_21.jpgScreen_Shot_2021-12-22_at_18.03.42.png
Message ID: 16522     Entry time: Fri Dec 17 19:19:42 2021     In reply to: 16519     Reply to this: 16525
Author: Koji 
Type: Update 
Category: SUS 
Subject: Remaining task for 2021 

I had the fear that any mistake in the electronics chain could have been the show stopper.

So I quickly checked the signal assignments for the ADC and DAC chains.

I had initial confusion (see below), but it was confirmed that the electronics chains (at least for LO1) are correct.

Note: One 70ft cable is left around the 1Y0 rack

 


There are a few points to be fixed:

- It looks like the ADC/DAC card # assignment has been messed up.

CDS ADC0 -> Cable label ADC1 -> AA A1 -> ...
CDS ADC1 -> Cable label ADC0 -> AA A0 -> ...
CDS DAC0 -> Cable label DAC2 -> AI D2 -> ...
CDS DAC1 -> Cable label DAC0 -> AI D0 -> ...
CDS DAC2 -> Cable label DAC1 -> AI D1 -> ...
(What is going on here... please confirm and correct as they become straight forward)

Once this puzzle was solved I could confirm reasonable connections from the end of the 70 cables to the ADC/DAC.

- We also want to change the ADC card assignment. The face OSEM readings must be assigned to ADC1 and the side OSEM readings to ADC0.
  My system wiring diagram needs to be fixed accordingly too.
  This is because the last channel of the first ADC (ADC0) is not available for us and is used for DuoTone.

Attachment 1: PXL_20211218_030145356.MP.jpg  4.886 MB  | Show | Show all
ELOG V3.1.3-